Here are the average rates of return for common stocks and Treasury bills for four different periods: 1900-1928 1929-1957 1958-1986 1987-2015 12,0% 4.9 9.8% 1.0 12.2% 6.1 12.0% 3.8 Stocks Treasury bills What was the risk premium on stocks for each of these periods?
Answer #1

Ans. Risk Premium = Estimated Return on Stocks - Estimated return on risk free bonds ( Treasury Bills )

Period :

1. 1900 - 1928 : 12% - 4.9 % = 7.1%

2. 1929 - 1957 : 9.8% - 1% = 8.8%

3. 1958 - 1986 : 12.2% - 6.1%. = 6.1%

4. 1987 - 2015 : 12% - 3.8% = 8.2%

Best of Luck !! !!